OpenStudy (anonymous):

The grades on the last math exam had a mean of 72%. Assume the population of grades on math exams is known to be distributed normally, with a standard deviation of 5%. Approximately what percent of students earn a score between 72% and 87%?

OpenStudy (anonymous):

49.9% 1% 50% 47.7%

OpenStudy (anonymous):

0.49865 or 49.9

OpenStudy (anonymous):

The Z value is 3. \[\frac{ 87-72 }{ 5 } = 3\] Look up the probability of Z = 3 at the table you should get 0.99865 Probability at 72 is 0.5 Since you are finding the probability in between 72 and 87: 0.99865-0.5=0.49865